The present age of Raman's mother is three times the present age Raman. After 5 years their ages will add to 66 years. Find their present ages
Answer by ankor@dixie-net.com(22740)

The present age of Raman's mother is three times the present age Raman.
m = 3r
After 5 years their ages will add to 66 years.
(m+5) + r+5) = 66
m + r + 10 = 66
m + r = 66 - 10
m + r = 56
replace m with 3r, from the 1st statement
3r + r = 56
4r = 56
r = 56/4
r = 14 yrs Raman's age
:
I'll let you find Mom's age.
:
Check the solutions in the statement
"After 5 years their ages will add to 66 years."